OpenStudy (allgemeine-klause01):

The point-slope form of the equation of a line that passes through points (8, 4) and (0, 2) is y – 4 = 1/4 (x – 8). What is the slope-intercept form of the equation for this line?

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

pooja195 (pooja195):

You are currently in the openstudy feedback section this belongs in the math section but no worries i can help ^_^ This is the equation in point slope form \[\huge~\rm~\bf~ y – 4 = 1/4 (x – 8) \] Slope intercept form is: \[\huge~\rm~\bf~y=mx+b \] To but the point slope equation into slope intercept we start by distributing the fraction first. 

pooja195 (pooja195):

Can you tell me what we have once we distribute? :)

OpenStudy (allgemeine-klause01):

i thought it was y=1/4x-4 but its wrong

OpenStudy (allgemeine-klause01):

and thx I'm new so learning where everything is so srry i post in wrong area

pooja195 (pooja195):

Ok let's do it step by step :) 1/4 * x= 1/4x 1/4 * -8=? No worries we all make mistakes ^_^

OpenStudy (allgemeine-klause01):

8/-32?

pooja195 (pooja195):

not quite :) \[\huge~\rm~\bf~\frac{ 1\times -8 }{ 4 \times 1 }=? \]

OpenStudy (allgemeine-klause01):

-8/4?

pooja195 (pooja195):

Good now divide -8 divded by 4=?

OpenStudy (allgemeine-klause01):

-2

pooja195 (pooja195):

Perfect :) \[\huge~\rm~\bf~y-4=1/4x-2\] Now simply add 4 to both sides -4+4=? -2+4=?

OpenStudy (allgemeine-klause01):

-8 and -6 (y=1/4x+2)??

pooja195 (pooja195):

The answer is correct y=1/4x+2 How ever -4+4=0 adding 4 to negative 4 cancels it out -2+4 makes it a postive 2

OpenStudy (allgemeine-klause01):

ok thx for clearing that up for me, and thx a lot for the help!

pooja195 (pooja195):

yw ^_^ Once again welcome to openstudy!

Zeronknight (zeronknight):

hello! I can see you are new here, having you considered reading the CoC? :) Also, glad you joined us! Just remember we are always here to help!

OpenStudy (allgemeine-klause01):

ok thx:)

Zeronknight (zeronknight):

@pooja195

pooja195 (pooja195):

:-)

eclipsedstar (eclipsedstar):

Don't forget to close your post when you feel like your question is adequately answered!

OpenStudy (allgemeine-klause01):

okay thx